# S2ST-Omni: An Efficient and Scalable Multilingual Speech-to-Speech Translation Framework via Seamless Speech-Text Alignment and Streaming Speech Generation

### v4

@@ 4.1.3 Baselines
Table 2: Offline S2ST performance of StreamSpeech, evaluated with BLASER 2.0.
[…]
BLASER 2.0 (Unsupervised)
UnitY 0.4467 0.4473 0.5090 0.5116
StreamSpeech 0.4486 0.4491 0.5155 0.5178
S2ST-Omni ??? ??? ??? ???
BLASER 2.0 (QE)
UnitY 3.1674 3.1772 3.3020 3.3278
StreamSpeech 3.1779 3.1872 3.3442 3.3669
S2ST-Omni ??? ??? ??? ???
BLASER 2.0 (Ref)
UnitY 3.1744 3.1965 3.2213 3.2638
StreamSpeech 3.1989 3.2200 3.3146 3.3525
S2ST-Omni ??? ??? ??? ???
@@ 4.2 Results and Analysis
To validate the effectiveness of the proposed approach, we compare the translation performance of S2ST-Omni against six state-of-the-art S2ST systems. A comprehensive summary of the results is provided in the Table 1 and Table 2.
@@ 5 CONCLUSIONS
In this study, we introduce S2ST-Omni, a unified and efficient framework for multilingual speech-to-speech translation. By decomposing the S2ST task into speech-to-text translation and text-to-speech synthesis, and integrating large-scale pre-trained models such as Whisper and Qwen 3.0 with a lightweight speech adapter, S2ST-Omni achieves high-quality translation while reducing reliance on parallel speech data. Furthermore, the adoption of a streaming speech generation model enables low-latency speech generation suitable for real-time applications. Extensive experiments on the CVSS benchmark confirm that S2ST-Omni achieves better translation quality than existing state-of-the-art systems, demonstrating its effectiveness and superiority.

### v5

@@ 3.2 Results and Analysis
To validate the effectiveness of the proposed approach, we compare the translation performance of S2ST-Omni against six state-of-the-art S2ST systems. A comprehensive summary of the results is provided in Table 1.
As shown in Table 1, S2ST-Omni consistently and significantly outperforms all baselines on the CVSS benchmark across the evaluated language pairs, with over 50% relative ASR-BLEU improvements in certain cases. These results highlight the strong generalization capability of the proposed framework across linguistically diverse settings. To be concrete, in the French $\rightarrow$ English direction, our S2ST-Omni achieves an ASR-BLEU score of 30.20, surpassing the strongest baseline (StreamSpeech, 28.45). This substantial improvement underscores the robustness of the propose method in handling phonetic variability and prosodic features characteristic of French. For German $\rightarrow$ English, the S2ST-Omni attains a 28.02 ASR-BLEU score, the highest among all language pairs as well, demonstrating its effectiveness in managing syntactic complexity and compound structures inherent to German. In the Spanish $\rightarrow$ English setting, S2ST-Omni achieves a remarkable 32.09 ASR-BLEU score, again outperforming all baselines by a wide margin. This result reflects the model’s ability to capture the rich morphological patterns of Spanish and generate fluent, accurate translations. In addition, despite the significant differences in phonology, syntax, and morphology among French, Spanish, and German, S2ST-Omni consistently delivers strong and comparable ASR-BLEU scores, further validating its stability and robustness across diverse linguistic scenarios.
@@ 4 CONCLUSIONS
In this study, we introduce an efficient and scalable framework, S2ST-Omni, for multilingual speech-to-speech translation. First, we propose an effective speech language model that integrates large-scale pre-trained multimodal models (Whisper and Qwen 3.0) with a lightweight speech adapter to achieve high-quality translation while reducing reliance on parallel speech data. Then, a two-stage progressive fine-tuning approach is presented to further improve the translation accuracy. Additionally, we employ a streaming speech generation model to balance the trade-off between the synthesis latency and translation quality. Extensive experiments on the CVSS benchmark confirm that S2ST-Omni achieves better translation quality than existing state-of-the-art systems, demonstrating its effectiveness and superiority.